OpenStudy (anonymous):

what is the equation for the line described. Give answers in standard form. (-5,4) m=-3/2 Can someone who knows how to do this please assist? Thank you.

OpenStudy (anonymous):

the point slope form of the line is y-4=(-3/2)(x+5)

OpenStudy (anonymous):

2(y-4)=-3(x+5) 2y=-3x-15+8 y=(-3x-7)/2

OpenStudy (anonymous):

thank you for the assistance